OpenStudy (anonymous):

Check all the ordered pairs that satisfy the equation below. y=2/3x A. (21,3/2) B. (21,14) C. (7,14) D. (28,42) E. (14,21) F. (42,28) *Help please

Mehek (mehek14):

plug their values into the equation

Mehek (mehek14):

\(\Huge{A}\\\dfrac{3}{2}=\dfrac{2}{3}*21\) solve that and tell me if it's true

OpenStudy (anonymous):

no ?

Mehek (mehek14):

move on to B then and solve

OpenStudy (anonymous):

I just know B is one of the answers .

Mehek (mehek14):

\(14=\dfrac{2}{3}*21\)

Mehek (mehek14):

ok so you have B as an answer now move onto C

Mehek (mehek14):

\(14=\dfrac{2}{3}*7?\)

OpenStudy (anonymous):

idk ?

Mehek (mehek14):

2*7 = 14 \(\dfrac{14}{3}=14?\)

OpenStudy (anonymous):

?

Mehek (mehek14):

does \(\dfrac{14}{3}\) equal to 14?

OpenStudy (anonymous):

nope

Mehek (mehek14):

so move onto D

Mehek (mehek14):

\(42=\dfrac{2}{3}*28?\)

OpenStudy (anonymous):

idk none of them , can you please give me the answers ?

Mehek (mehek14):

nope I have to help you understand how to solve

Mehek (mehek14):

what is \(28*2?\)

OpenStudy (anonymous):

ok :( 56

Mehek (mehek14):

\(\dfrac{56}{3}=?\)

OpenStudy (anonymous):

18?

Mehek (mehek14):

so not D

OpenStudy (anonymous):

k next?

Mehek (mehek14):

\(21=\dfrac{2}{3}*14?\)

OpenStudy (anonymous):

9 ?

Mehek (mehek14):

so not E either

Mehek (mehek14):

\(28=\dfrac{2}{3}*42?\)

OpenStudy (anonymous):

so what are the right ones ? I know B Is pne

Mehek (mehek14):

we're checking the last answer choice the others were wrong except for B

Mehek (mehek14):

so solve for F and then we'll see

OpenStudy (anonymous):

okay 28

Mehek (mehek14):

so B and F are our answers

OpenStudy (anonymous):

okay thanks /